OpenStudy (anonymous):

18 5/12+8 10/17+23 7/12+28+7/17

OpenStudy (anonymous):

@arjund12

OpenStudy (anonymous):

no

OpenStudy (lizz123):

The answer is 49.

OpenStudy (anonymous):

answer is 79 not 49. correct answer is 79 :)

OpenStudy (xxdragoonxx):

10/17+8+18+23+28+7/17+7/12+5/12 8+18+23+28=77 = 10/17+77+7/17+7/12+5/12 1*7*17+1*5*17+1*2^2*3*7+1*2^2*3*10+2^2*3*17*77/204 10/17+77+7/17+7/12+5/12 = 77/1+7/17+7/12+5/12+10/17 12 = 2^2*3 12 = 2^2*3 1*2^2*3*17 = 204 = 2^2*3*17*77/204+1*7*17/204+1*5*17/204+1*2^2*3*7/204+1*2^2*3*7/204+1*2^2*3*10/204 = 1*7*17+1*5*17+1*2^2*3*7+1*2^2*3*10+2^2*3*17*77/204 = 2^2*3978+204/204 = 204(1+78)/204 = 79*204/204 = 79

OpenStudy (anonymous):

i agree with @xxDragoonxx. exactly :)